# Contrastive Localized Language-Image Pre-Training

### v1

@@ 3.3 CLOC Model Architecture
We implement the promptable embedding introduced in Section 3.2 with minimal extra modules on top of the original CLIP image and text encoders. As illustrated in Figure 2, the original CLIP model remains the same for computing ${\mathcal{L}}_{\text{CLIP}}$. For computing ${\mathcal{L}}_{\text{{CLOC}}}$/${\mathcal{L}}_{\text{grounding}}$, the image embedding is re-used from the CLIP ViT but before the pooled projection and normalization $f_{I}^{\prime}$. To extract the region embedding ${\bm{z}}=\texttt{Prompter}({\bm{l}},f_{I}^{\prime}({\bm{x}}))$ from the image, we consider the location representation ${\bm{l}}$ as two coordinates (top-left and bottom-right corners of a box), each vectorized by positional encoding. The Prompter is a simple and lightweight one-layer transformer encoder. It takes the positional encodings prepended with the sequence of image tokens from ViT together as the input, and outputs the region embedding with a pooled projection layer. For the grounding loss, we re-use the same CLIP text encoder for encoding the region captions ${\bm{z}}=\texttt{Prompter}(f_{T}({\bm{y}}),f_{I}^{\prime}({\bm{x}}))$ to predict the bounding boxes with a two-layer MLP head. Overall, CLOC only adds additional learnable parameters of the lightweight Prompter. Note, that the main overheads in a single forward are from encoding the image via the ViT – CLOC reuses it for multiple prompts.
@@ 3.4 Discussions on Design Choices and Extensions
To train our model with ${\mathcal{L}}_{\textsc{CLOC}}$ in Equation 4, it requires extracting region embeddings from the image features given the bounding boxes. A perhaps straightforward alternative could be Region-of-Interest (RoI) pooling/alignment (He et al., 2017) from the spatial image feature of ViT before pooling. RoI operations are popular, especially in the object detection literature. [^1] We observe a withdrawn arXiv preprint (https://arxiv.org/abs/2401.06397) proposes to extract RoI features for CLIP regional contrastive learning but only focus on dense vision tasks, not MLLMs. However, as will be evidenced by worse performance in Section 5, we found it suboptimal for CLOC pre-training here for several reasons. First, unlike object detection datasets that typically contain golden labels, here the pseudo-labels are much noisier on the large-scale web-crawled images. Therefore, the resulting RoI features may be inaccurate due to the imprecise bounding boxes, making model training less effective. Second, unlike dense vision tasks that directly rely on the spatial features, MLLM has a transformer decoder that consists of several attention layers such that the constraint of semantics in the spatial feature space becomes somewhat indirect. Our Prompter mimics such inductive bias in pre-training via a single-attention-layer encoder that may leverage better global context reasoning compared to RoIs.
@@
While region annotations introduce location information, a concern of contrastive learning on the regional captions may be that there are many similar objects within an image (e.g., “boats” in the harbor) or a mini-batch. To mitigate such concerns, we apply two tricks. First, fortunately, we found it sufficient to sample a few regions per image for each update, e.g., we set $M=4$ in Equation 2 in experiments. Second, we can filter similar texts when computing the negatives in the contrastive loss. More specifically, we ignore the pairs of $\big{(}{\bm{z}}_{i}^{(m)},f_{T}({\bm{y}}_{j}^{(m^{\prime})})\big{)}$ in the denominators of both ${\mathcal{L}}_{R\rightarrow T/{T\rightarrow R}}$, if $\texttt{sim}\big{(}f_{T}({\bm{y}}_{i}^{(m)}),f_{T}({\bm{y}}_{j}^{(m^{\prime})})\big{)}>0.9$, without gradients on $f_{T}$.
4 Visually-Enriched and Spatially-Localized Captioning Pipeline
Figure 3: Overview of the Visually-Enriched and Spatially-Localized (VESL) captioning pipeline. We leverage an existing open-vocabulary detector (e.g., OWLv2) that typically predicts bounding boxes on the images, and assigns the labels from the given text phrase candidates. Previous methods do not tailor how the text phrases are prepared and often use the alt-text attached to the images, which is prone to insufficient region descriptions. We found it crucial for CLOC to train on data from our VESL that re-captions images with the visually-enriched captioner VeCap (Lai et al., 2024) for better visual concept exploitation of the detector.
Table 1: Region-text dataset statistics. We summarize the text token length for both images and regions. Partial statistics of the proprietary datasets revealed by their papers. The $20$M subset of GRIT is released at: https://huggingface.co/datasets/zzliang/GRIT; we removed the invalid images.
Dataset # of images regions per image image caption length region text length
@@ 4 Visually-Enriched and Spatially-Localized Captioning Pipeline
GRIT (proprietary) (Peng et al., 2023) 91M 1.5 4.7
GRIT (released, clean) (Peng et al., 2023) 17M 1.8 17.2 4.6
Florence-2 (proprietary) (Xiao et al., 2024) 126M 5.4 70.5 2.6
OWLv2 (proprietary) (Minderer et al., 2024) 2B
WiT labeled w/ Minderer et al. (2024) 300M 5.1 17.1 3.9
@@
As discussed in Section 1 and 3.1, a key bottleneck of CLOC training is the region-text annotation datasets in terms of both the data size scales and the label quality, since there are no public datasets with region-text annotations at scales large enough for contrastive pre-training.
Inspired by recent works that re-caption images with visually-enriched captions for better CLIP training, we make a step further for Visually-Enriched and Spatially-Localized (VESL) labeler which generates more fine-grained captions at the region level. The goal of VESL is, given an image (possibly with the original web-crawled alt-text), annotate it with the grounded bounding boxes each associated with a caption in natural language for optimizing Equation 2 in Section 3.2.
Concretely, VESL is constructed as a pseudo-labeling pipeline with the following steps:
  1. 1.
    Image re-captioning with visual concept exploitation: We follow the VeCap framework (Lai et al., 2024) to generate long, diverse, and detailed image captions.
  2. 2.
    Region phrase candidates extraction: Inspired by Zhang et al. (2022), we apply name entity recognition (NER) to extract leaf entities from the visually-enriched captions as potential candidate phrases describing a region inside the image.
  3. 3.
    Open-vocabulary detection with extracted phrases: we generate the final region-text annotations via a pre-trained open-vocabulary detector queried with the phrases extracted from Step 2 to match the bounding boxes proposed by the detector. We adopted the OWLv2 detector (Minderer et al., 2024) which contains the CLIP image/text encoder with the detection head. The boxes with detection confidence larger than $0.1$ are kept as the region location and the phrases matched with the highest score are considered as their captions.
@@
Our pre-training data consists of two parts: ($i$) image-text pairs, and ($ii$) region-text pairs. For image-text pairs, we reproduce the image re-captioning pipeline from VeCap (Lai et al., 2024), and generate synthetic captions for WiT-300M (Wu et al., 2024) and DFN-5B (Fang et al., 2023) images. For region-text pairs, we pseudo-label WiT-300M and a 2B-image subset of DFN-5B using our VESL pipeline. In VESL, we adopted the official OWLv2 L/14 model (Minderer et al., 2024) as the open-vocabulary detector. [^2] OWLv2 CLIP L/14 ST+FT in: https://github.com/google-research/scenic/tree/main/scenic/projects/owl_vit All images are pseudo-labeled with $448\times 448$ resolution, where a maximum number of $20$ phrase queries are sampled for moderate computation budget. Table 1 summarizes the statistics of existing region-text datasets and ours. Notably, we also ablate annotating WiT-300M following Minderer et al. (2024) and found it detects less objects with longer region text, likely due to verbose $n$-grams of alt-text are in lower quality than our approach, as discussed in the remarks. Examples and pseudo codes are in Appendix B.
5 Experiments
5.1 Setup
Pre-training.
We follow OpenAI-CLIP (Radford et al., 2021) to train both our CLIP baseline model and CLOC model using a similar budget of around 14B images seen [^3] To avoid confusion, we will refer to CLIP as training with the standard CLIP loss on the same data as CLOC and refer to OpenAI-CLIP as the public pre-trained checkpoint released from Radford et al. (2021).. For a fair comparison, we use the same hyper-parameters and images for both the CLIP baseline and CLOC. We experimented with the ViT B/16 and L/14 architectures, pre-trained with $224\times 224$ and $336\times 336$ image resolutions, respectively. All parameters are trained end-to-end from scratch. We implement the codebase based on open-source AXLearn framework [^4] https://github.com/apple/axlearn in JAX (Bradbury et al., 2018). We provide hyper-parameters and more details in Appendix A.
Evaluation tasks.
We evaluate our image encoders across a wide range of downstream tasks. First, we assess performance on ImageNet image classification (Deng et al., 2009; Shankar et al., 2020) and COCO retrieval (Lin et al., 2014). Second, we construct region-level tasks, including COCO object recognition and region-text retrieval using the GRIT dataset (Peng et al., 2023). Furthermore, we show CLOC is particularly useful for MLLMs, validated by the Ferret model (You et al., 2023) which requires fine-grained image understanding for referring and grounding tasks. We also evaluate on general multimodal benchmarks using LLaVA-1.5 (Liu et al., 2023) and LLaVA-NeXT (Liu et al., 2024) [^5] We use the codebase: https://github.com/xiaoachen98/Open-LLaVA-NeXT., which both use the 7B Vicuna LLM. For all evaluation tasks, we use the same official hyper-parameters, fine-tuning datasets, and codebase for all the image encoders we experimented with, without specific tuning. More details are provided in each subsection and in Appendix A.
5.2 Image and Region Classification and Retrieval Tasks
The proposed CLOC training framework enables the encoder to produce not only image embedding but also region embeddings. It can directly be used for region-level tasks without further training, in analogy to the zero-shot capability of CLIP on images. To evaluate such capability and also for fast development and ablation study, we first construct several region-level zero-shot tasks.
Besides image-level evaluation like ImageNet classification and COCO image-text retrieval, we additionally construct region-level tasks, including region object recognition and region-text retrieval. More specifically, the region-level tasks leverage the labeled bounding boxes in the evaluation set for CLOC to extract region embedding. For region retrieval, we use a validation set of the GRIT dataset (Peng et al., 2023) and encode both the image regions and the region captions. For region classification, the class names are encoded as text embedding ($80$ / $1203$ classes for COCO / LVIS, respectively), and the highest cosine similarity for each region embedding is predicted as its class.
@@ 5.2 Image and Region Classification and Retrieval Tasks
Table 2: Zero-shot evaluation on image-level tasks (accuracy of ImageNet (IN) classification, recall@1 of COCO retrieval) and region-level tasks (mAcc of region object recognition on COCO and LVIS, recall@10 of GRIT region retrieval), using ViT-B/16 as the default encoder backbone. The indentation with different symbols denotes removing (–) or changing a component ($\circ$).
Models Training data Image tasks Region tasks Avg.
Image Region COCO (i2t) COCO (t2i) INv1 INv2 GRIT (r2t) GRIT (t2r) COCO Recog. LVIS Recog. Image Region
@@
We highlight important variables for the performance in Table 2 with the following observations:
  • CLOC performs decently on region-level tasks [^6] For reference, in a different data setup, Wu et al. (2023) reports 46.5% mAcc on the same COCO region classification task, trained using $320\times 320$ COCO training images. In contrast, our approach achieves over 70% mAcc, pre-trained on a $224\times 224$ large-scale web-crawled dataset with object annotations (thus might not be fair comparisons). while maintaining strong performance on image-level metrics ( 2 vs.  8 13).
  • The Prompter is an important ingredient for CLOC’s success to go beyond CLIP without a compromise ( 3 vs.  4; 8 vs.  9; 13 vs.  14). We replace the Prompter with RoI alignment to extract region features and train with ${\mathcal{L}}_{\text{{CLOC}}}$. We found it performs much worse on region-level tasks than CLOC, possibly due to such strong constraints of RoI features being difficult to learn on the noisy labels with the CLIP loss as discussed in Section 3.4.
  • VESL outperforms Minderer et al. (2024) baseline approach, as the visually-enriched captions improve image retrieval tasks (as expected (Lai et al., 2024)), while also offering versatile visual concepts as text candidates for the OV detector, supporting Section 4 ( 3 vs.  5; 13 vs.  15).
  • Given the same captions in VESL, OWLv2 slightly outperforms the GLIPv2 detector ( 3 vs.  6).
@@
  • Region tasks work well when sampling $2$ or $4$ boxes per image, making CLOC practical ( 12).
  • Scaling up region labels seems saturated at 300M images on region tasks ( 3 8 13), while we found it will further improve in MLLM tasks as will be shown in Table 3.
@@
Overall, CLOC not only achieves strong performance on image-level tasks, but unlocks a new capability for zero-shot region-level tasks. We have validated our design choices for architectures, training, and data. Below, the complete setup 13 will be used as default if not specified.
Table 3: Results on Ferret-Bench for referring and grounding VQA, based on Ferret (You et al., 2023) equipped with different image encoders. Models are evaluated with OpenAI gpt-4o API instead of the deprecated gpt-4-0314 in the paper. replace Ferret visual sampler with Prompter; see Section 5.3 for details.
Method ViT Region Alignment # of images w/ region labels Referring Description Referring Reasoning Grounding in Conversation Avg. ( $\Delta$ to CLIP)
@@
OpenAI-CLIP L/14 None None 50.8 55.4 45.7 50.6
@@
5.3 Referring and Grounding with Ferret
We evaluate CLOC by replacing the CLIP ViT encoder with our CLOC ViT as a drop-in replacement. We follow the official codebase [^7] We use the official Ferret codebase: https://github.com/apple/ml-ferret. for training the Ferret model. We further consider a variant based on Ferret: the Ferret model implements a spatial-aware visual sampler that samples image features from the region specified in the question. We replace the sophisticated visual sampler with our simple Prompter introduced in Section 3.3 to extract region embedding with ${\bm{z}}=\texttt{Prompter}({\bm{l}},f_{I}^{\prime}({\bm{x}}))$ instead, as illustrated in Figure 1(right).
Table 4: Results on referred object classification (LVIS), referring expression comprehension ($0.5$ IoU on RefCOCO, RefCOCO+, RefCOCOg), and phrase grounding ($0.5$ IoU on Flickr30k Entities) with Ferret. Shikra: baseline in Chen et al. (2023). Ferret: replace visual sampler with CLOC prompter.
Model Encoder LVIS RefCOCO RefCOCO+ RefCOCOg Flickr Avg.
@@ 5.3 Referring and Grounding with Ferret
Shikra OpenAI-CLIP L/14 57.8 67.7 n/a 87.0 90.6 80.2 81.6 87.4 72.1 82.3 82.2 75.8 76.5 -
Ferret OpenAI-CLIP L/14 79.4 67.9 69.8 87.5 91.4 82.5 80.8 87.4 73.1 83.9 84.8 80.4 82.2 80.8
@@
Table 5: Results on multimodal benchmarks using LLaVA-1.5/NeXT with ViT-L/14 and Vicuna-7B.
@@
CLIP LLaVA-1.5 59.3 53.3 62.2 30.0 86.7 1451.4 254.3
CLOC LLaVA-1.5 64.3 54.9 62.7 31.5 87.3 1482.0 288.9
CLIP Open-LLaVA-NeXT 67.3 61.4 63.5 38.5 87.9 1486.1 279.6
CLOC Open-LLaVA-NeXT 69.5 61.9 64.2 40.2 88.3 1451.1 312.5

### v2

@@ 3.3 CLOC Model Architecture
We implement the promptable embedding introduced in Section 3.2 with minimal extra modules on top of the original CLIP. As illustrated in Figure 2, the CLIP model remains the same for computing ${\mathcal{L}}_{\text{CLIP}}$. For ${\mathcal{L}}_{\text{{CLOC}}}$/${\mathcal{L}}_{\text{grounding}}$, the image embedding is re-used from the CLIP ViT but before the pooled projection and normalization $f_{I}^{\prime}$. To extract the region embedding ${\bm{z}}=\texttt{Prompter}({\bm{l}},f_{I}^{\prime}({\bm{x}}))$ from the image, we consider the location representation ${\bm{l}}$ as two coordinates (top-left and bottom-right corners of a box), each vectorized by positional encoding. The Prompter is a lightweight one-layer transformer encoder. It takes the positional encodings prepended with the image tokens from ViT together as the input, and outputs the region embedding with a pooled projection layer. For the grounding loss, we re-use the same CLIP text encoder for encoding the region captions ${\bm{z}}=\texttt{Prompter}(f_{T}({\bm{y}}),f_{I}^{\prime}({\bm{x}}))$ to predict the bounding boxes with a two-layer MLP head. Overall, CLOC only adds lightweight additional parameters of the heads. Note that, the main overheads during forward are from ViT image encoding – CLOC reuses it for multiple prompts.
Figure 3: Our Visually-Enriched and Spatially-Localized (VESL) captioning pipeline. We leverage an existing open-vocabulary detector (e.g., OWLv2) that predicts bounding boxes on the images and assigns the labels from the given text phrase candidates. Previous methods often use the alt-text attached to the images, which is prone to insufficient region descriptions. We found it crucial to re-caption images with the visually-enriched captioner VeCap (Lai et al., 2024) for better visual concept exploitation for the detector.
@@ 3.4 Discussions on Design Choices and Extensions
Training with ${\mathcal{L}}_{\textsc{CLOC}}$ in Equation 4 requires extracting region embeddings from the image features given the bounding boxes. Another alternative could be Region-of-Interest (RoI) pooling/alignment (He et al., 2017) from the spatial image feature of ViT before pooling. RoI operations are popular, especially in the object detection literature. However, as will be evidenced by worse performance in Section 5, we found spatial pooling an over-strong premise for CLOC pre-training here for several reasons.
First, unlike object detection datasets that typically contain golden labels, here the pseudo-labels are much noisier on the large-scale web-crawled images. The resulting RoI features may be inaccurate due to the imprecise bounding boxes, making model training less effective. Second, unlike dense vision tasks that directly rely on the spatial features, MLLM has a transformer decoder that consists of several attention layers such that the constraint of semantics in the spatial feature space becomes somewhat indirect. Our Prompter mimics such inductive bias in pre-training via a single-attention-layer encoder that may leverage better global context reasoning compared to RoIs.
@@
While region annotations introduce location information, a concern of contrastive learning may be similar objects within an image (e.g., “boats” in the harbor) or a mini-batch. To mitigate it, we apply two tricks. First, fortunately, we found it sufficient to sample a few regions per image for each update, e.g., we set $M=4$ in Equation 2 in experiments. Second, we can filter similar texts when computing the negatives in the contrastive loss. More specifically, we ignore the pairs of $\big{(}{\bm{z}}_{i}^{(m)},f_{T}({\bm{y}}_{j}^{(m^{\prime})})\big{)}$ in the denominators of both ${\mathcal{L}}_{R\rightarrow T/{T\rightarrow R}}$, if $\texttt{sim}\big{(}f_{T}({\bm{y}}_{i}^{(m)}),f_{T}({\bm{y}}_{j}^{(m^{\prime})})\big{)}>0.9$, without gradients on $f_{T}$.
4 VESL Captioning Pipeline
Table 1: Region-text dataset statistics. We summarize the text token length for both images and regions. Partial statistics of the proprietary datasets revealed by their papers. The $20$M subset of GRIT is released; we removed the invalid URLs.
Dataset # of images regions per image image text len. region text len.
@@ 4 VESL Captioning Pipeline
GRIT (prop.) (Peng et al., 2023) 91M 1.5 4.7
GRIT (released) (Peng et al., 2023) 17M 1.8 17.2 4.6
Florence-2 (prop.) (Xiao et al., 2024) 126M 5.4 70.5 2.6
OWLv2 (prop.) (Minderer et al., 2024) 2B
WiT labeled w/ (Minderer et al., 2024) 300M 5.1 17.1 3.9
@@
As discussed in Section 1 and 3.1, a key bottleneck of CLOC is the region-text datasets in terms of both the data size and the label quality, since there are no public datasets with region-text annotations at scales large enough for contrastive pre-training. Inspired by recent works that enrich image captions for better CLIP training, we make a step further for Visually-Enriched and Spatially-Localized (VESL) captioning that generates fine-grained captions at region level. The goal of VESL is, given an image with the original web-crawled alt-text, annotates it with the grounded bounding boxes each associated with a caption in natural language for optimizing Equation 2 in Section 3.2.
Concretely, VESL is a pseudo-labeling pipeline with the following steps, with pseudo codes in Appendix C:
1. Re-captioning with visual concept exploitation: We follow the VeCap2 (Lai et al., 2024, 2025) to generate long, diverse, and detailed image captions.
2. Region phrase candidates extraction: We apply name entity recognition (NER) to extract phrases from the visually-enriched captions as candidate describing a region inside the image, inspired by Zhang et al. (2022).
3. Open-vocabulary detection with extracted phrases: the final region-text annotations are generated by a pre-trained open-vocabulary detector. It matches the phrases extracted from Step 2 to the bounding boxes proposed by the detector. We adopted OWLv2 detector (Minderer et al., 2024) which combines the CLIP image/text encoders with detection heads. The boxes with confidence larger than $0.1$ are kept as the region location and the most confident phrases are considered as their captions.
@@
Our pre-training data consists of two parts: ($i$) image-text pairs, and ($ii$) region-text pairs. For image-text pairs, we reproduce the image re-captioning pipeline from VeCap (Lai et al., 2024), and generate synthetic captions for WiT-300M (Wu et al., 2024) and DFN-5B (Fang et al., 2023) images. For region-text pairs, we pseudo-label WiT-300M and a 2B-image subset of DFN-5B using our VESL pipeline. In VESL, we adopted the official OWLv2 L/14 model (Minderer et al., 2024) as the open-vocabulary detector. All images are pseudo-labeled with $448\times 448$ resolution, where a maximum number of $20$ phrase queries are sampled for moderate computation budget. Table 1 summarizes the statistics of existing region-text datasets and ours. Notably, we also ablate annotating WiT-300M following (Minderer et al., 2024) and found it detects less objects with longer region text, likely due to verbose $n$-grams of alt-text are in lower quality than our approach, as discussed in the remarks.
@@
Table 2: Zero-shot evaluation on image-level tasks (recall@1 of COCO retrieval and accuracy of ImageNet (IN) classification) and region-level tasks (recall@10 of GRIT region retrieval and mAcc of region object recognition on COCO and LVIS), using ViT-B/16 as the default encoder backbone. The indentation with different symbols denotes removing (–) or changing a component ($\circ$).
Models Training Labels Image tasks Region tasks Avg.
Image Region COCO-i2t COCO-t2i INv1 INv2 GRIT-r2t GRIT-t2r RecCOCO RecLVIS Image Region
@@
5 Experiments
5.1 Setup (more details in Appendix B)
Pre-training.
We follow OpenAI-CLIP (Radford et al., 2021) to train both our CLIP baseline model and CLOC model using a similar budget of around 14B images seen. For a fair comparison, we use the same hyper-parameters and images for both the CLIP baseline and CLOC. We experimented with the ViT B/16 and L/14 architectures, pre-trained with $224\times 224$ and $336\times 336$ image resolutions, respectively. All parameters are trained end-to-end from scratch. We implement the in JAX (Bradbury et al., 2018).
Evaluation tasks.
The image encoders are evaluated across a wide range of downstream tasks. First, we assess performance on ImageNet image classification (Deng et al., 2009; Shankar et al., 2020) and COCO retrieval (Lin et al., 2014). Second, we construct region-level tasks, including COCO object recognition and region-text retrieval using the GRIT dataset (Peng et al., 2023). Furthermore, we show CLOC is particularly useful for MLLMs, validated by the Ferret model (You et al., 2023) which requires fine-grained image understanding for referring and grounding tasks. We also evaluate on general multimodal benchmarks using LLaVA-1.5 (Liu et al., 2023) and Open-LLaVA-NeXT (Liu et al., 2024; Chen & Xing, 2024), which both use the 7B Vicuna LLM. For all evaluation tasks, we use the same official hyper-parameters, fine-tuning datasets, and codebase for all the image encoders we experimented with, without specific tuning.
5.2 Image and Region Classification and Retrieval
CLOC encoder produces not only image embedding but also region embeddings. It can be used directly for region-level tasks without further training, in analogy to the zero-shot capability of CLIP on images. This emergent capability enables us to construct region-level zero-shot tasks for fast development and ablation study.
In addition to image-level evaluation like ImageNet classification and COCO image-text retrieval, we additionally construct region-level tasks, including region object recognition and text retrieval. More specifically, the region-level tasks leverage the labeled bounding boxes in the evaluation set for CLOC to extract region embedding. For region retrieval, we use a validation set of the GRIT dataset (Peng et al., 2023) and encode both the image regions and the region captions. For region classification, the class names are encoded as text embedding ($80$ / $1203$ classes for COCO / LVIS, respectively), and the highest cosine similarity for each region embedding is predicted as its class. We highlight important variables for the performance in Table 2 with the following observations:
  • CLOC performs decently on region-level tasks [^1] For reference, in a different setup, Wu et al. (2023) reports 46.5% mAcc on the COCO region classification task, trained with $320\times 320$ COCO images directly. In contrast, our approach achieves over 70% mAcc, pre-trained on a $224\times 224$ web-crawled dataset with our object labels (thus not a fair comparisons). with strong image-level performance ( 2 vs.  8 13).
  • The Prompter is an important ingredient for CLOC’s success to go beyond CLIP ( 3 8 13 vs.  4 9 14). We replace the Prompter with RoI alignment to extract region features and train with ${\mathcal{L}}_{\text{{CLOC}}}$ (similar to (Shi et al., 2025)). We found it performs much worse on region-level tasks, possibly due to difficulties of strong RoI constraints and the noisy labels as discussed in Section 3.4.
  • VESL helps, as the visually-enriched captions improve image retrieval tasks (as expected (Lai et al., 2024)) and the versatile visual concepts candidates facilitate the OV detector, supporting Section 4 ( 3 13 vs.  5 15).
  • OV detector OWLv2 $>$ GLIPv2 in VESL ( 3 vs.  6).
@@ 5.2 Image and Region Classification and Retrieval
  • Practically, sampling $2$/$4$ boxes works ( 12) well already.
  • Scaling up images saturated on region tasks but further improved on MLLM tasks ( 3 8Table 3.
@@
Overall, CLOC not only achieves strong performance on image-level tasks, but unlocks zero-shot region-level capability. Below, with our design choices validated, 13 will be used by default if not specified.
Table 3: Ferret-Bench for referring and grounding VQA, based on Ferret (You et al., 2023) equipped with different image encoders. Models are evaluated with OpenAI gpt-4o API. replace Ferret visual sampler with Prompter; Details in Section 5.3.
Method ViT Region Align #images w/ region labels Ref-Descript. Ref-Reason. Ground Conv. Avg.
$\Delta$CLIP
@@
OAI-CLIP L/14 None None 50.8 55.4 45.7 50.6
@@
Table 4: Results on referred-LVIS object classification, referring expression comprehension ($0.5$ IoU on RefCOCO, RefCOCO+, RefCOCOg), and phrase grounding ($0.5$ IoU on Flickr30k Entities). Ferret : replace visual sampler in Ferret with CLOC prompter.
Model Encoder RefLVIS RefCOCO RefCOCO+ RefCOCOg Flickr Avg.
@@
Shikra (Chen et al., 2023) OpenAI-CLIP L/14 57.8 67.7 n/a 87.0 90.6 80.2 81.6 87.4 72.1 82.3 82.2 75.8 76.5 -
Ferret (Zhang et al., 2024) OpenAI-CLIP L/14 79.4 67.9 69.8 87.5 91.4 82.5 80.8 87.4 73.1 83.9 84.8 80.4 82.2 80.8
@@
Table 5: Results on multimodal benchmarks using LLaVA-1.5 and Open-LLaVA-NeXT with ViT-L/14 and Vicuna-7B.
@@
LLaVA-1.5
CLIP 59.3 53.3 62.2 30.0 86.7 1451.4 254.3
CLOC 64.3 54.9 62.7 31.5 87.3 1482.0 288.9
Open-LLaVA-NeXT
CLIP 67.3 61.4 63.5 38.5 87.9 1486.1 279.6
CLOC 69.5 61.9 64.2 40.2 88.3 1451.1 312.5
5.3 Referring and Grounding with Ferret
We evaluate CLOC by replacing the CLIP ViT encoder with our CLOC ViT as a drop-in replacement. We use the official codebase for training the Ferret model. We further consider a variant based on Ferret: the Ferret model implements a spatial-aware visual sampler that samples image features from the region specified in the question. We replace the sophisticated visual sampler with our simple Prompter introduced in Section 3.3 to extract region embedding with ${\bm{z}}=\texttt{Prompter}({\bm{l}},f_{I}^{\prime}({\bm{x}}))$ instead, as illustrated in Figure 1(right).